# Praxis-VLM: Vision-Grounded Decision Making via Text-Driven Reinforcement Learning

### v2

@@ Abstract
Vision Language Models exhibited immense potential for embodied AI, yet they often lack the sophisticated situational reasoning required for complex decision-making. This paper shows that VLMs can achieve surprisingly strong decision-making performance when visual scenes are represented merely as text-only descriptions, suggesting foundational reasoning can be effectively learned from language. Motivated by this insight, we propose Praxis-VLM, a reasoning VLM for vision-grounded decision-making. Praxis-VLM employs the GRPO algorithm on textual scenarios to instill robust reasoning capabilities, where models learn to evaluate actions and their consequences. These reasoning skills, acquired purely from text, successfully transfer to multimodal inference with visual inputs, significantly reducing reliance on scarce paired image-text training data. Experiments across diverse decision-making benchmarks demonstrate that Praxis-VLM substantially outperforms standard supervised fine-tuning, exhibiting superior performance and generalizability. Further analysis confirms that our models engage in explicit and effective reasoning, underpinning their enhanced performance and adaptability.
@@ 1 Introduction
Meanwhile, advancements in large language models (LLMs), such as DeepSeek-R1 [10] and OpenAI o1 [11], highlight the potential of multi-step reasoning for tackling complicated tasks. Recent efforts have attempted to enhance VLMs with sophisticated reasoning capability from text-based models [12, 13, 14]. These approaches typically utilize reasoning-oriented LLMs to generate long-form reasoning chains, which are then used to supervise VLM training. However, they rely heavily on large-scale, high-quality vision data paired with textual ground-truth answers, which are notoriously expensive and laborious to curate across diverse real-world scenarios [15, 16, 17]. The challenge of obtaining such paired image-text training data becomes even more pronounced in real-world embodied decision-making contexts with diverse situations.
@@
This data acquisition challenge consequently motivates us to investigate the fundamental nature of decision-making abilities and their reliance on direct multimodal training. An essential question then arises: Is the core of decision-making ability exclusively tied to direct multimodal experience? If not, there may be more cost-effective pathways to improve the multimodal decision-making ability of VLMs. To address this, we conduct a preliminary analysis (§ 2) and find a surprisingly effective yet underexplored alternative: when vision-grounded situations are represented by textual descriptions, even standard VLMs could achieve comparative or even improved performance on complex multimodal decision-making benchmarks like VIVA [18] and PCA-Bench [19]. This observation sparked our central hypothesis: fundamental decision-making knowledge and reasoning capability can be learned primarily from language-based representations and then effectively transferred to visually-situated contexts during inference. This notion resonates with the mental model theory [20], which posits that humans construct internal, often language-based, representations of situations to reason, predict outcomes, and guide their decisions, later applying these models to sensory experiences.
@@
Motivated by this insight, we conduct a systematic investigation into this hypothesis and propose Praxis-VLM, a reasoning VLM that learns high-level decision-making principles from language and applies this “praxis” within vision-grounded environments. Specifically, we begin by constructing a text-based training corpus where visual situations are articulated through descriptive language, mitigating the need for image-text paired data. Then, to foster robust and transferable reasoning—the ability to “think before decide”—we employ a Reinforcement Learning (RL) approach. Specifically, we employ GRPO algorithm [21] with multi-stage training to encourage the model to generate explicit reasoning chains before reaching a decision. To facilitate efficient learning, we introduce a novel adaptive R1 reward that targets different skills at each training stage. This learned capacity for deliberate reasoning then transfers effectively when Praxis-VLM encounters and processes actual visual inputs during multimodal inference. Illustrative examples are shown in Figure 1.
@@
In summary, our work makes three-fold contributions: (1) We show the potential of leveraging language as a primary medium for instilling transferable reasoning skills in VLMs for situational decision-making; (2) Building on this, we propose a text-based reinforcement learning paradigm and introduce Praxis-VLM, a novel model that learns high-level decision-making principles from language and grounds them in concrete, multimodal scenarios via an adaptive R1 reward; (3) Through extensive experiments and analyses across three diverse benchmarks, we demonstrate Praxis-VLM’s superior reasoning and generalization capabilities for embodied decision-making, charting a cost-efficient path for scalable VLM reasoning.
@@ 2 Preliminary Analysis
To investigate alternative pathways for developing decision-making skills , we preliminarily analyze the performance of VLMs under different input conditions. We evaluated Qwen2.5-VL [25], on two vision-grounded decision-making benchmarks: VIVA [18] and PCA-Bench [19]. They require the model to choose the best action from multiple options based on an image depicting a specific situation. We compare two settings: (1) using the original image as input situation, and (2) using a textual description of the situation, either a caption generated by GPT-4o or taken from the dataset’s annotations, in place of the image.
@@
The results, presented in Figure 2, reveal a compelling insight: the text situation setting could demonstrate performance comparable to, or even slightly better than, the VLM operating directly on the image-text paired input. This observation suggests that the fundamental reasoning and decision logic required for navigating these human-centric and embodied-agent scenarios can be substantially captured and learned from symbolic, textual representations alone.
Such findings resonate with human cognitive development, where abstract knowledge, reasoning skills, and decision-making strategies are often acquired through language, detached from an immediate sensory perception of the real situation. Based on this insight, we hypothesize that VLMs can similarly benefit from acquiring reasoning capabilities primarily through text-based learning. Therefore, our core methodological premise is to cultivate sophisticated reasoning and decision-making policies using rich, text-only situational descriptions paired with desired outcomes. The ultimate aim is to transfer these textually-learned reasoning skills effectively to multimodal inference for the model to ground its decisions in the actual visual information. It enables the model to leverage the text data for reasoning development while retaining the VLM’s ability to perceive and act in vision-grounded contexts.
@@ 3 Method: Learning to Reason and Decide from Text
Figure 3: Overview of Praxis-VLM: Learning transferable reasoning from text-only data for multimodal decision-making. The framework involves creating synthetic text-based training data, training the VLM with GRPO with adaptive rewards on this data to develop reasoning, and applying the learned reasoning to make decisions in vision-grounded situations during inference.
Our primary goal is to enhance the reasoning and decision-making capabilities of VLMs in vision-grounded scenarios. Recognizing the challenges in acquiring large-scale image-text pairs for training and inspired by §$\S$ 2, we propose a novel paradigm to learn decision-making skills from text-only data and transfer these skills to multimodal inference. Here, we employ Reinforcement Learning to foster the model’s ability to generate explicit reasoning processes for complex decision-making. As illustrated in Figure 3, our framework involves three key phases: (1) Creating a text-based decision-making dataset, (2) Optimizing the VLM’s reasoning and decision-making using GRPO [21], and (3) Deploying the enhanced VLM for decision-making with actual visual input during inference.
@@ 3.1 Problem Formulation and Model Setup
Inspired by §$\S$ 2, importantly, our training strategy focuses on enhancing reasoning capabilities primarily through text. Therefore, during the training phase, visual inputs $x^{S}$ are replaced by their textual descriptions $x^{S}_{text}$, and only the language model components of the VLM are updated. This leverages the scalability of text data for knowledge acquisition. Yet, during the inference phase, the entire trained VLM architecture, including the vision encoder, is used to process the image-text input pair $(x^{S},x^{Q})$, allowing the textually-learned reasoning skills to be applied to visual situations.
@@ 3.2 Text-Based Decision-Making Data Construction
Leveraging recent advances in LLM-based data synthesis [17, 27], we employ GPT-4o [28] for generation. Our process begins with 10 manually crafted seed questions serving as in-context examples. To further maximize the diversity of the situations, we implement a batch generation approach where GPT-4o produces 10 samples per time, followed by a deduplication step. This strategy allows an effective generation of varied scenarios and questions, yielding a dataset comprising 10K training samples and 1K validation samples. More details of data creation are in Appendix A.3.
@@ 3.4 Multi-Stage GRPO Training with Adaptive R1 Reward
For this decision-making RL training, the adaptive rewards aim to encourage both comprehensive thinking and accurate decisions. We use a combination of: (1) $R_{\text{format}}$: Continues to ensure adherence to the thinking-answering structure; (2) $R_{\text{accuracy}}$: Rewards the correctness of the answer; (3) $R_{\text{len}}$: Encourages the model to generate deliberate and longer reasoning chains. Our observations indicate that $R_{\text{len}}$ is effective in promoting a more comprehensive consideration of the situation and action candidates. Contrary to some recent work suggesting that long reasoning chains might be redundant [33, 34], our results show that encouraging longer reasoning can lead to more thorough situational analysis for more complex situations. We will provide detailed discussions on this in §$\S$ 5.2. This adaptive reward strategy across stages enables efficient training by targeting different skills, including format adherence, logical computation, and complex decision-making, sequentially.
@@ 4.1 Tasks and Datasets
We consider VIVA and PCA-Bench as in-domain benchmarks, as they align with the typical image-text input and decision-making formats our model is trained on. In contrast, EgoNormia serves as an out-of-domain benchmark, introducing additional challenges such as sequential and temporal reasoning over video frames and egocentric perception. These datasets offer a rigorous and diverse testbed for evaluation. We follow the original data splits and prompts provided by each benchmark.
@@ 4.2 Baselines and Implementations
Implementation Details. For both GRPO and SFT training, we finetune full model parameters. For GRPO training, we set rollout N to 5 and KL divergence coefficient to 0.01. During inference, we leverage VLLM Library [36] with greedy decoding. More details are in Appendix A.2.
@@ 5.4 Exploring Praxis-VLM’s Reasoning: What Does It Consider?
To gain deeper insights into the reasoning capabilities learned by Praxis-VLM, we analyze its generated reasons. We first prompt GPT-4o to generate keyphrases that summarize the reasoning aspects of each sample. These keyphrases are then clustered across the dataset. As shown in Figure 5, this analysis reveals four primary dimensions characterizing the model’s reasoning process.
@@ 5.5 Understanding Failures: Error Analysis Through the Lens of Reasoning
Figure 6: Illustrative examples of common failure cases for Praxis-VLM, identified by analyzing the reasoning generated for incorrect predictions. Complete outputs are in Appendix B.
These failure modes highlight that while encouraging explicit reasoning is beneficial, challenges remain in ensuring deep and accurate contextual understanding, robust prioritization under complex constraints, and reliable alignment with nuanced social and ethical norms. Addressing these aspects within the reasoning process is a key direction for future research. More samples are in Appendix B.
@@ 6 Related Work
VLMs in Embodied Decision-Making. The quest to enable intelligent agents to make informed decisions in real-world, situated environments is central to embodied AI research [2, 22, 37, 38]. VLMs have emerged as a powerful foundation for such agents, demonstrating significant promise in applications like robotics, autonomous navigation, and interactive task planning [39, 40, 41, 42]. These models integrate visual perception with language understanding to interpret and interact with their surroundings. However, a persistent challenge lies in equipping VLMs with the capacity for multi-step reasoning required to navigate and act effectively in nuanced and dynamic situations. Our work focuses on enhancing such sophisticated reasoning abilities crucial for robust decision-making.
Reasoning in VLMs. The ability to reason is fundamental to effective decision-making. Traditional methods often rely on fine-tuning VLMs on large multimodal datasets tailored to specific tasks or reasoning styles [16, 43, 44]. More recently, approaches have emerged that encourage VLMs to generate explicit reasoning steps, often leveraging RL to optimize performance on complex tasks [13, 35, 45, 46, 47]. While these RL-based methods have shown success in improving reasoning, they typically necessitate extensive training on datasets comprised of paired image-text data. Our text-driven RL diverges from this by proposing a more data-efficient pathway to instill reasoning.
@@ 7 Conclusion
Discussions on Limitations and Future Work
@@
Ethics Statement and Broader Impacts
Appendix A Experiment Details
A.1 Data Statistics
@@ A.2 Implementation Details
A.3 Details on Text-Based Data Creation
Figure 8: An example of a synthetic data sample used for our text-based training. It comprises a textual situation description, a multiple-choice decision-making question, and the target answer. The reasoning chain, also generated by GPT-4o, is used for Reason SFT training, and not utilized during our GRPO training phase.
@@ A.3 Details on Text-Based Data Creation
Appendix B Sample Outputs

### v3

@@ Abstract
Vision Language Models exhibit impressive performance for various tasks, yet they often lack the sophisticated situational reasoning required for complex decision-making. This paper shows that VLMs can achieve surprisingly strong decision-making performance when visual scenes are replaced by textual descriptions, suggesting foundational reasoning can be effectively learned from language. Motivated by this insight, we propose Praxis-VLM, a reasoning VLM for vision-grounded decision-making. Praxis-VLM employs the GRPO algorithm on textual scenarios to instill robust reasoning capabilities, where models learn to evaluate actions and their consequences. These reasoning skills, acquired purely from text, successfully transfer to multimodal inference with visual inputs, significantly reducing reliance on scarce paired image-text training data. Experiments across diverse decision-making benchmarks demonstrate that Praxis-VLM substantially outperforms standard supervised fine-tuning, exhibiting superior performance and generalizability. Further analysis confirms that our models engage in explicit and effective reasoning, underpinning their enhanced performance and adaptability.
@@ 1 Introduction
Meanwhile, advancements in recent large language models (LLMs) highlight the potential of multi-step reasoning for tackling complicated tasks [10, 11]. Recent efforts have attempted to enhance VLMs with sophisticated reasoning capability from text-based models [12, 13, 14]. These approaches typically utilize reasoning-oriented LLMs to generate reasoning chains, which are then used to supervise VLM training. However, they rely heavily on large-scale, high-quality vision data paired with textual ground-truth answers, which are notoriously expensive and laborious to curate across diverse real-world scenarios [15, 16, 17]. The challenge of obtaining such paired image-text training data becomes even more pronounced in real-world decision-making contexts with diverse situations.
@@
This data acquisition challenge consequently motivates us to investigate the fundamental nature of decision-making abilities and their reliance on direct multimodal training. An essential question then arises: Is the core of decision-making ability exclusively tied to direct multimodal experience? If not, there may be more cost-effective pathways to improve the multimodal decision-making ability of VLMs. To address this, we conduct a preliminary analysis (§ 2) and find a surprisingly effective yet underexplored alternative: when visual situations are represented by textual descriptions, even standard VLMs could achieve comparative or even improved performance on complex multimodal decision-making benchmarks like VIVA [18] and PCA-Bench [19]. This observation sparked our central hypothesis: fundamental decision-making and reasoning capabilities can be disentangled from visual perception and learned primarily through language-based representations, which can then be effectively transferred to visually grounded contexts during inference. This notion resonates with the mental model theory [20], which posits that humans construct internal, often language-based, representations of situations to reason, predict outcomes, and guide their decisions, later applying these internal models to sensory experiences and act upon real-world situations.
@@
Motivated by this insight, we propose Praxis-VLM, a reasoning VLM that learns high-level decision-making principles from language and applies this “praxis” to vision-grounded environments. Specifically, we begin by constructing a text-based training corpus where visual situations are articulated through text descriptions, mitigating the need for image-text paired data. Then, to foster robust and transferable reasoning—the ability to “think before decide”—we employ a Reinforcement Learning (RL) approach. Specifically, we employ GRPO algorithm [21] with multi-stage training to encourage the model to generate explicit reasoning chains before reaching a decision. To facilitate effective learning, we introduce a novel adaptive R1 reward that targets different skills at each training stage. The reasoning abilities acquired through this process are then transferred when Praxis-VLM processes actual visual inputs during multimodal inference. Illustrative examples are shown in Figure 1.
@@
In summary, our main contributions are: (1) We show the potential of leveraging language as a medium for instilling transferable reasoning skills in VLMs for situational decision-making; (2) Building on this, we propose a text-based RL paradigm and introduce Praxis-VLM, a novel model that learns high-level decision-making principles from language and grounds them in concrete, multimodal scenarios via an adaptive R1 reward; (3) Through extensive experiments and analyses across three diverse benchmarks, we demonstrate Praxis-VLM’s superior reasoning and generalization capabilities for embodied decision-making, charting a practical and data-efficient path for VLM training.
@@ 2 Preliminary Analysis
To investigate alternative pathways for developing decision-making skills , we preliminarily analyze the performance of VLMs under different input conditions. We evaluated Qwen2.5-VL [25], on two vision-grounded decision-making benchmarks: VIVA [18] and PCA-Bench [19]. Both benchmarks are framed as Visual Question Answering tasks, requiring the model to choose the best action from multiple options based on an image depicting a specific situation. We compare two settings: (1) using the original image as input situation, and (2) using a textual description of the situation, either a caption generated by GPT-4o or taken from the dataset’s annotations, in place of the image.
@@
The results, presented in Figure 2, reveal a compelling insight: the text situation setting shows performance comparable to, or even slightly better than, the VLM operating directly on the image input. This observation suggests that the fundamental reasoning and decision logic required for navigating these human-centric and embodied-agent scenarios can be substantially captured and learned from textual representations alone.
Such findings resonate with human cognitive development, where abstract knowledge, reasoning skills, and decision-making strategies are often acquired through language, detached from an immediate perception of the real situation [26]. Based on this insight, we hypothesize that VLMs can similarly benefit from acquiring reasoning capabilities primarily through text-based learning. Therefore, our core methodological premise is to cultivate sophisticated reasoning and decision-making policies using rich, text-only situational descriptions paired with desired outcomes. The ultimate aim is to transfer these textually-learned reasoning skills effectively to multimodal inference for the model to ground its decisions in the actual visual information. It enables the model to leverage the text data for reasoning development while retaining the VLM’s ability to perceive and act in vision-grounded contexts.
@@ 3 Method: Learning to Reason and Decide from Text
Figure 3: Overview of Praxis-VLM: Learning transferable reasoning from text-only data for multimodal decision-making. The process involves (1) constructing synthetic text-based training data where situations are represented through textual descriptions, (2) training the VLM on this data using RL with adaptive rewards to develop reasoning skills, and (3) transferring the learned reasoning to vision-grounded decision-making tasks during inference.
Our primary goal is to enhance the reasoning and decision-making capabilities of VLMs in vision-grounded scenarios. Recognizing the challenges in acquiring large-scale image-text pairs for training and inspired by $\S$ 2, we propose a novel paradigm to learn decision-making skills from text-only data and transfer these skills to multimodal inference. Here, we employ Reinforcement Learning to foster the model’s ability to generate explicit reasoning processes for complex decision-making. As illustrated in Figure 3, our framework involves three key phases: (1) Creating a text-based decision-making dataset, (2) Optimizing the VLM’s reasoning and decision-making using GRPO [21], and (3) Deploying the enhanced VLM for decision-making with actual visual input during inference.
@@ 3.1 Problem Formulation and Model Setup
Inspired by $\S$ 2, importantly, our training strategy focuses on enhancing reasoning capabilities primarily through text. Therefore, during the training phase, we synthesize text-based data $\mathcal{D}=\{(x^{S}_{text},x^{Q},y)\}$, where the visual inputs $x^{S}$ are replaced with textual descriptions $x^{S}_{text}$. The model is then trained on this data by updating only the language model components of the VLM, leveraging the scalability of text for efficient knowledge acquisition. Yet, during the inference phase, the entire trained VLM architecture, including the vision encoder, is used to process the image-text input pair $(x^{S},x^{Q})$, allowing the textually-learned reasoning skills to be applied to visual situations.
@@ 3.2 Text-Based Decision-Making Data Construction
Leveraging recent advances in LLM-based data synthesis [17, 28], we employ GPT-4o [29] for data construction. Specifically, we first craft 10 seed questions as in-context examples to guide data generation, then prompt GPT-4o to produce additional samples. To maximize data diversity, we adopt a batch generation approach, producing 10 samples at a time, followed by deduplication. This strategy allows an effective generation of varied scenarios and questions, yielding a final dataset of 10K training samples and 1K validation samples. Importantly, our method requires no manual filtering or intensive curation, enabling fast, domain-agnostic training data generation while reducing reliance on costly image-pair datasets. More details of data creation are in Appendix B.3.
@@ 3.4 Multi-Stage GRPO Training with Adaptive R1 Reward
For this decision-making RL training, the adaptive rewards aim to encourage both comprehensive thinking and accurate decisions. We use a combination of: (1) $R_{\text{format}}$: Continues to ensure adherence to the thinking-answering structure; (2) $R_{\text{accuracy}}$: Rewards the correctness of the answer; (3) $R_{\text{len}}$: Encourages the model to generate deliberate and longer reasoning chains. Our observations indicate that $R_{\text{len}}$ is effective in promoting a more comprehensive consideration of the situation and action candidates. Contrary to some recent work suggesting that long reasoning chains might be redundant [34, 35], our results show that encouraging longer reasoning can lead to more thorough situational analysis for more complex situations. We will provide detailed discussions on this in $\S$ 5.2. This adaptive reward strategy across stages enables efficient training by targeting different skills, including format adherence, logical computation, and complex decision-making, sequentially.
@@ 4.1 Tasks and Datasets
All tasks are formalized as VAQ and we employ accuracy as the evaluation metric. We consider VIVA and PCA-Bench as in-domain benchmarks, as they align with the typical image-text input and decision-making formats our model is trained on. In contrast, EgoNormia serves as an out-of-domain benchmark, introducing additional challenges such as sequential and temporal reasoning over video frames and egocentric perception. These datasets offer a rigorous and diverse testbed for evaluation. We follow the original data splits and prompts provided by each benchmark.
@@ 4.2 Baselines and Implementations
Implementation Details. For both GRPO and SFT training, we finetune full model parameters. For GRPO training, we set rollout N to 5 and KL divergence coefficient to 0.01. During inference, we leverage VLLM Library [37] with greedy decoding. More details are in Appendix B.2.
@@ 5.4 Exploring Praxis-VLM’s Reasoning: What Does It Consider?
To gain deeper insights into the reasoning ability learned by Praxis-VLM, we analyze its generated reasons. We first prompt GPT-4o to generate keyphrases that summarize the reasoning aspects of each sample. These keyphrases are then clustered across the dataset. As shown in Figure 5, this analysis reveals four primary dimensions characterizing the model’s reasoning process.
@@ 5.5 Understanding Failures: Error Analysis Through the Lens of Reasoning
Figure 6: Illustrative examples of common failure cases for Praxis-VLM, identified by analyzing the reasoning generated for incorrect predictions. Complete outputs are in Appendix C.
These failure modes highlight that while encouraging explicit reasoning is beneficial, challenges remain in ensuring deep and accurate contextual understanding, robust prioritization under complex constraints, and reliable alignment with nuanced social and ethical norms. Addressing these aspects within the reasoning process is a key direction for future research. More samples are in Appendix C.
@@ 6 Related Work
VLMs in Situational Decision-Making. The quest to enable intelligent agents to make informed decisions in real-world, situated environments is central to embodied AI research [2, 22, 38, 39]. VLMs have emerged as a powerful foundation for such agents, demonstrating significant promise in applications like robotics, autonomous navigation, and interactive task planning [40, 41, 42, 43]. These models integrate visual perception with language understanding to interpret and interact with their surroundings. However, a persistent challenge lies in equipping VLMs with the capacity for multi-step reasoning required to navigate and act effectively in nuanced and dynamic situations. Our work focuses on enhancing such sophisticated reasoning abilities crucial for robust decision-making.
Reasoning in VLMs. The ability to reason is fundamental to effective decision-making. Recent studies have explored VLMs’ visual reasoning capabilities across various tasks, including visual question answering and commonsense reasoning [44, 45, 46, 47]. Traditional methods often rely on fine-tuning VLMs on large multimodal datasets tailored to specific tasks or reasoning styles [16, 48, 49]. More recently, approaches have emerged that encourage VLMs to generate explicit reasoning steps, often leveraging RL to optimize performance on complex tasks [13, 36, 50, 51, 52]. While these RL-based methods have shown success in improving reasoning, they typically necessitate extensive training on datasets comprised of paired image-text data. Our text-driven RL diverges from this by proposing a more data-efficient pathway to instill reasoning.
@@ 7 Conclusion
Ethics Statement and Broader Impacts
Acknowledgement
This work is supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25200821), the Innovation and Technology Fund (Project No. PRP/047/22FX), and PolyU Internal Fund from RC-DSAI (Project No. 1-CE1E).
Appendix A Discussions on Limitations and Future Work
@@
Appendix B Experiment Details
B.1 Data Statistics
@@ B.2 Implementation Details
B.3 Details on Text-Based Data Creation
Figure 8: Left: An example of a synthetic data sample used for our text-based training. It comprises a textual situation description, a multiple-choice decision-making question, and the target answer. Right: The reasoning chain, also generated by GPT-4o, is used for Reason SFT training, and not utilized during our GRPO training phase.
@@ B.3 Details on Text-Based Data Creation
Data Diversity Analysis. To assess the diversity of the generated scenarios, we conduct analysis by prompting GPT-4o to cluster the generated textual situations into topical categories. The clustering results indicate a broad coverage across varied scenarios: (1) Workplace Performance and Personal Issues, (2) Resource Allocation, (3) Project Management, (4) Balancing Competing Interests, (5) Policy, Rules & Enforcement, (6) Ethical Dilemmas, (7) Interpersonal Conflict, (8) Emergency Handling, (9) Navigating Setbacks, (10) Event Planning & Logistics, and (11) Balancing Inclusivity and Majority Preference.
Appendix C Sample Outputs